# Reactive Transport Modeling with Physics-Informed Machine Learning for Critical Minerals Applications

### v2

@@ 1. INTRODUCTION AND MOTIVATION
Understanding mass transport is crucial for predicting plume behavior when flow, diffusion, and reactions occur. In many cases, chemical reactions are sufficiently fast (e.g., in situ leaching scenarios), allowing them to be approximated as instantaneous [Dentz et al., 2011]. In this study, flow through porous media, chemical transport, and fast bimolecular reactions are treated as interconnected processes: the flow field informs transport and reaction dynamics, and all three components together determine reactive transport behavior. The modeling framework proceeds in three coupled stages: first solving the flow problem to obtain the velocity field, then computing the dispersion tensor for transport, and finally solving the steady-state diffusion-reaction equation to model fast chemical reactions.
@@
Recent studies have used machine learning to accelerate reactive transport simulations by replacing chemical solvers with trained surrogates or reduced models [Leal et al., 2020; Kyas et al., 2022]. While effective within operator-splitting frameworks, these approaches treat transport and reactions separately. In contrast, PINNs approach used in this work integrates flow, transport, and reactions in a coupled framework. Hence, PINNs can resolve all the governing physics directly, by solving the reactive transport problem monolithically and provide a flexible and well-suited framework for inverse modeling and data-scarce scenarios. Physics-informed machine learning (PIML) has been applied to porous media flow [Fraces and Tchelepi, 2021; Mao et al., 2020; Wang et al., 2020], from subsurface energy systems [Latrach et al., 2024] to vascular cooling [Jagtap et al., 2023]. Though termed PIML or PINNs in the literature [Harp et al., 2021; Jagtap et al., 2023], both share the same principles; we adopt “PINNs” throughout this paper.
@@ 2.1. PINNs modeling framework
If $\mathbf{x}$ represents spatial points within the domain, let $n$ denote the total number of collocation points. The continuous coordinates in the domain are referred to as spatial points, while the discrete subset used to evaluate the PINNs residuals is referred to as collocation points. The input layer receives a total of $n$ collocation points, each with spatial coordinates ($x$, $y$), forming an input tensor of shape ($n$, 2) for two dimensional (2D) problems, which are then passed through the hidden layers. At each neuron, an activation function is applied, and the output variables are computed as the initialized weights and biases propagate through the specified hidden layers ($l$) and neurons/nodes ($n_{l}$). Mathematically, the operation inside each neuron can be expressed as follows:
@@ 3. VERIFICATION OF FLOW SUBPROBLEM
For all three cases, the boundary value equation presented in Eq. (2.9) is used to train the PINNs model and solve for pressure and velocity at each spatial point. The solution is then tested and validated against the finite element method (FEM) solution obtained utilizing fluid flow models—second-order Lagrange elements on a triangular mesh with a standard continuous Galerkin formulation—available in COMSOL Multiphysics [2018].
@@ 3.3. Inclined patch
All the results combined together in Figs.  68 complete the three patch tests, each highlighting different scenarios of permeability heterogeneity. The developed PINNs code successfully captures the heterogeneity in all cases, demonstrating its effectiveness for such flow problems. The results closely match those obtained using FEM, confirming the accuracy of the approach. Thus, this section provides a validation of the PINNs’ capability to solve these types of flow problems, hence answering the first scientific question.
@@ 4.2. Problem setup
$\mathrm{D}=\left[{\begin{array}[]{cc}\cos({\theta})&\sin({\theta})\\ -\sin({\theta})&\cos({\theta})\\ \end{array}}\right]\;\left[{\begin{array}[]{cc}k_{1}&0\\ 0&k_{2}\\ \end{array}}\right]\;\left[{\begin{array}[]{cc}\cos({\theta})&-\sin({\theta})\\ \sin({\theta})&\cos({\theta})\\ \end{array}}\right]$
where $\theta=\pi/6$, $k_{1}=10000$ and $k_{2}=1$ for extreme anisotropy. With this configuration, the governing equations in Eqs.  (2.13)–(2.15) are solved over the square plate with a hole.
@@ 4.3. Numerical results
On the other hand, PINNs produced solutions that adhered to the maximum principle without requiring additional constraints or modifications to the loss function or model architecture, responding to the second science question. This highlights the advantage of PINNs in solving diffusion problems while strictly maintaining physical principles, making them a powerful and efficient tool for such applications.
@@ 5.1. Uniform flow velocity
A square domain with $L_{x}=L_{y}=1$ is chosen. At the left boundary, a unit pressure is prescribed while the middle third right boundary is prescribed with zero pressure. The remaining boundaries are prescribed with zero velocity as shown in Fig.  11 A). The body force $\mathbf{b(x)}$ is assumed to be zero. The coefficient of permeability is taken as $k=1$ over the domain. Flow subproblem is solved utilizing the governing equations in Eqs.  (2.9)- (2.12) solving for the pressure profile as presented in the Fig.  11 B) with the maximum pressure at the left boundary and minimum at the middle third of right boundary, satisfying the boundary conditions.
@@ 5.2. Non-uniform explicit velocity
After evaluating the performance of PINNs on a fast bimolecular reaction with uniform flow velocity, the study is extended to examine the case of random and non-uniform flow velocity. To achieve this, an explicit velocity is selected referring from Mudunuru and Nakshatrala [2016], with its $x$ and $y$ components, $\mathbf{v}(x)$ and $\mathbf{v}(y)$, defined as:
@@
The velocity field obtained is used to compute the diffusivity tensor using Eq.  (2.20). The invariants $\Psi_{A}$ and $\Psi_{B}$ represented in Eqs. (2.26) and  (2.28) are solved using PINNs and plotted in Fig. 14 B) and C) respectively. Similar to the flow velocity case, the values of invariants $\Psi_{A}$ and $\Psi_{B}$ range from 0 to 1. $\Psi_{A}$ corresponds to the concentration of reactant A and product C, exhibiting maximum diffusivity in the upper half of the domain where reactant A is dominant, as per the problem setup. Similarly, $\Psi_{B}$ represents reactant B and product C, exhibiting peak diffusion in the lower half, where reactant B is introduced in the problem setup.
@@
Finally, the concentration of reactants A and B along with product C are computed using the invariants $\Psi_{A}$ and $\Psi_{B}$ and presented in Fig. 15 . The mathematical calculation of the concentration is done using Eqs. (2.35)–(2.37). Reactant A exhibits its highest concentration at the top left boundary, gradually decreasing toward the right boundary. Similarly, reactant B reaches its maximum concentration at the bottom left boundary and diminishes as it moves rightward. The formation of product C is observed at the interaction zone of reactants A and B, displaying a striking wavy concentration pattern. Product C begins forming at the left boundary and develops into a distinct plume extending toward the right boundary. Notably, the plume shifts slightly upward within the domain, which aligns with expectations, given that the coefficient $n_{B}$ is twice that of $n_{A}$, facilitating the upward movement of the plume.
@@ 6. CLOSURE
We have developed a physics-informed modeling framework for fast bimolecular reactions, incorporating tests for flow and reactive transport in porous media. In applications where fluid flow uncertainties and the risk of chemical contamination are concerns, PINNs—a novel machine learning framework—offers a powerful solution. Unlike traditional data-driven methods, PINNs is trained using the governing physics rather than relying on large datasets.
This approach has broad applicability, ranging from critical minerals to subsurface energy systems such as geothermal energy and geological carbon storage, where fluid flow and chemical transport play critical roles. A key challenge in such subsurface applications is the scarcity of sufficient data. PINNs address this limitation by leveraging the system of PDEs and boundary conditions, enabling accurate modeling without the need for extensive experimental or field data. PINNs is trained by minimizing the total loss, which is a combination of individual losses calculated based on how well the solution at each collocation point satisfies the given PDE and boundary conditions. The system of equations varies depending on the nature of the problem—for instance, whether it involves flow or transport, as in this study. During the development of the PINNs framework and solving the problems, several key observations were made, which are summarized below:
@@
Data supporting the findings of this study are available from the corresponding authors (KBN) on a reasonable request. The codes used in this paper will be made available and hosted at EMSL-Computing GitHub (https://github.com/EMSL-Computing/) upon publication.

### v3

@@ 1. INTRODUCTION AND MOTIVATION
Understanding mass transport is crucial for predicting plume behavior when flow, diffusion, and reactions occur. In many cases, chemical reactions are sufficiently fast (e.g., in situ leaching scenarios), allowing them to be approximated as instantaneous [Dentz et al., 2011]. Mixing-limited reactions in porous media arise when reaction kinetics are fast relative to transport, such that reaction rates are controlled by incomplete mixing driven by diffusion and dispersion. This behavior has been extensively studied using theoretical, numerical, and experimental approaches, highlighting the strong dependence of reactive plume evolution on pore-scale heterogeneity and mixing dynamics rather than intrinsic reaction rates. Previous studies have employed pore-scale simulations, upscaled continuum models, and particle- and volume-averaged approaches to quantify mixing-controlled reaction rates and effective reaction terms [Valocchi et al., 2019]. However, accurately capturing sharp reaction fronts and dispersive mixing at the continuum scale remains challenging due to numerical diffusion, scale dependence, and the need for fine spatial resolution. Comprehensive reviews of these challenges and modeling strategies are provided by. In this study, flow through porous media, chemical transport, and fast bimolecular reactions are treated as interconnected processes: the flow field informs transport and reaction dynamics, and all three components together determine reactive transport behavior. The modeling framework proceeds in three coupled stages: first solving the flow problem to obtain the velocity field, then computing the dispersion tensor for transport, and finally solving the steady-state diffusion-reaction equation to model fast chemical reactions.
@@
Recent studies have used machine learning to accelerate reactive transport simulations by replacing chemical solvers with trained surrogates or reduced models [Leal et al., 2020; Kyas et al., 2022]. While effective within operator-splitting frameworks, these approaches treat transport and reactions separately. In contrast, PINNs approach used in this work integrates flow, transport, and reactions in a coupled framework. Hence, PINNs can resolve governing physics that are enforced directly within each subproblem using PINNs, and the overall reactive transport process is solved in a sequential, physics-consistent manner, where the flow solution provides inputs to the reactive transport model and provide a flexible and well-suited framework for inverse modeling and data-scarce scenarios. Physics-informed machine learning (PIML) has been applied to porous media flow [Fraces and Tchelepi, 2021; Mao et al., 2020; Wang et al., 2020], from subsurface energy systems [Latrach et al., 2024] to vascular cooling [Jagtap et al., 2023]. Though termed PIML or PINNs in the literature [Harp et al., 2021; Jagtap et al., 2023], both share the same principles; we adopt “PINNs” throughout this paper.
@@ 2.1. PINNs modeling framework
If $\mathbf{x}$ represents spatial points within the domain, let $n$ denote the total number of collocation points. The continuous coordinates in the domain are referred to as spatial points, while the discrete subset used to evaluate the PINNs residuals is referred to as collocation points. The input layer receives a total of $n$ collocation points, each with spatial coordinates ($x$, $y$), forming an input tensor of shape ($n$, 2) for two dimensional (2D) problems. For each collocation point, a random value—generated using the default random value generator in the JAX library— is assigned as the first step of the training process, which is then passed through the NN structure. At each neuron, an activation function is applied, and the output variables are computed as the initialized weights and biases propagate through the specified hidden layers ($l$) and neurons/nodes ($n_{l}$). Mathematically, the operation inside each neuron can be expressed as follows:
@@ 2.4.2. Applicability of fast bimolecular reactions to critical mineral science
A summary of workflow steps is as follows:
  1. (1)
    Input & setup: Specify domain/boundary conditions, heterogeneous permeability $K(\mathbf{x})$, porosity $\phi$, fluid properties, stoichiometry $(n_{A},n_{B},n_{C})$, and dispersion parameters $(\alpha_{L},\alpha_{T},D_{m})$.
  2. (2)
    Flow PINN (Darcy/mixed form): Train a PINN to solve for pressure $p(\mathbf{x})$ and Darcy velocity $\mathbf{v}(\mathbf{x})$ using a PDE-residual loss plus boundary-condition losses; validate with patch tests. The trained velocity field $\mathbf{v}(\mathbf{x})$ is then fixed and passed to the transport step.
  3. (3)
    Compute anisotropic dispersion tensor: Construct
    $\mathbf{D}(\mathbf{x})\;=\;D_{m}\mathbf{I}\;+\;\mathbf{D}_{\mathrm{disp}}\!\big(\mathbf{v}(\mathbf{x}),\alpha_{L},\alpha_{T}\big),$
    and, if used, the effective tensor $\mathbf{D}(\mathbf{x})/\phi$.
  4. (4)
    Transport–reaction via invariants (reaction part): Instead of training a network directly on stiff reaction source terms, solve two reaction-invariant transport equations for $\Psi_{A}(\mathbf{x},t)$ and $\Psi_{B}(\mathbf{x},t)$. These invariants are constructed so that the fast bimolecular reaction cancels from the governing equations, leaving uncoupled diffusion equations (implemented here through anisotropic diffusion with $\mathbf{D}(\mathbf{x})$).
  5. (5)
    Algebraic fast-reaction closure (reconstruction): Recover species concentrations from invariants using stoichiometry and non-negativity/mutual-exclusion rules, e.g.,
    $c_{A}=\max\!\Big(\frac{\Psi_{A}}{n_{A}}-\frac{\Psi_{B}}{n_{B}},0\Big),\qquad c_{B}=\max\!\Big(\frac{\Psi_{B}}{n_{B}}-\frac{\Psi_{A}}{n_{A}},0\Big),\qquad c_{C}=\min\!\Big(\frac{\Psi_{A}}{n_{A}},\frac{\Psi_{B}}{n_{B}}\Big),$
    (or equivalent piecewise forms). This step concerns the “ML implementation for the reaction” in our approach: the network learns the invariant transport fields, while the reaction itself is enforced via an explicit, physics-based algebraic projection that is consistent with the instantaneous-reaction limit.
  6. (6)
    Outputs & checks: Generate $c_{A},c_{B},c_{C}$ fields, verify non-negativity, and compare plume metrics (e.g., front location and product evolution) against reference solutions.
@@ 3. VERIFICATION OF FLOW SUBPROBLEM
For all three cases, the boundary value equation presented in Eq. (2.9) is used to train the PINNs model and solve for pressure and velocity at each spatial point, a total of $100\times 100$ for each case. The solution is then tested and validated against the finite element method (FEM) solution obtained utilizing fluid flow models—second-order Lagrange elements on a triangular mesh with a standard continuous Galerkin formulation—available in COMSOL Multiphysics [2018].
@@ 3.3. Inclined patch
All the results combined together in Figs.  68 complete the three patch tests, each highlighting different scenarios of permeability heterogeneity. The error analysis and convergence plots are presented in the supplementary material to provide robustness in the solution obtained. The developed PINNs code successfully captures the heterogeneity in all cases, demonstrating its effectiveness for such flow problems. The results closely match those obtained using FEM, confirming the accuracy of the approach. Thus, this section provides a validation of the PINNs’ capability to solve these types of flow problems, hence answering the first scientific question.
@@ 4.2. Problem setup
$\mathrm{D}=\left[{\begin{array}[]{cc}\cos({\theta})&\sin({\theta})\\ -\sin({\theta})&\cos({\theta})\\ \end{array}}\right]\;\left[{\begin{array}[]{cc}\lambda_{1}&0\\ 0&\lambda_{2}\\ \end{array}}\right]\;\left[{\begin{array}[]{cc}\cos({\theta})&-\sin({\theta})\\ \sin({\theta})&\cos({\theta})\\ \end{array}}\right]$
where $\theta=\pi/6$, the principal components of the anisotropic diffusion $\lambda_{1}=10000$ and $\lambda_{2}=1$ for extreme anisotropy. With this configuration, the governing equations in Eqs.  (2.13)–(2.15) are solved over the square plate with a hole.
@@ 4.3. Numerical results
On the other hand, PINNs produced solutions that adhered to the maximum principle without requiring additional constraints or modifications to the loss function or model architecture, responding to the second science question. Unlike classical discretization methods, PINNs approximate the solution as a smooth global function and enforce the governing equations pointwise at collocation points via the loss function. This strong-form enforcement, together with the absence of element-wise polynomial interpolation, mitigates spurious oscillations commonly associated with Gibbs phenomena. Consequently, non-physical negative concentrations are implicitly penalized during training, allowing the PINNs solution to adhere to the maximum principle without additional constraints. This highlights the advantage of PINNs in solving diffusion problems while strictly maintaining physical principles, making them a powerful and efficient tool for such applications.
@@ 5.1. Uniform flow velocity
A square domain with $L_{x}=L_{y}=1$ is chosen. At the left boundary, a unit pressure is prescribed while the middle third right boundary is prescribed with zero pressure. The remaining boundaries are prescribed with zero velocity as shown in Fig.  11 A). The body force $\mathbf{b(x)}$ is assumed to be zero. The coefficient of permeability is taken as $k=1$ over the domain. Flow subproblem is solved utilizing the governing equations in Eqs.  (2.9)–(2.12) solving for the pressure profile as presented in the Fig.  11 B) with the maximum pressure at the left boundary and minimum at the middle third of right boundary, satisfying the boundary conditions.
@@ 5.2. Non-uniform explicit velocity
After evaluating the performance of PINNs on a fast bimolecular reaction with uniform flow velocity, the study is extended to examine the case of random and non-uniform flow velocity. To achieve this, an explicit velocity is explicitly prescribed and is selected referring from Mudunuru and Nakshatrala [2016], with its $x$ and $y$ components, $\mathbf{v}(x)$ and $\mathbf{v}(y)$, defined as:
@@
The velocity field obtained is used to compute the diffusivity tensor using Eq.  (2.20). The invariants $\Psi_{A}$ and $\Psi_{B}$ represented in Eqs. (2.26) and  (2.28) are solved using PINNs and plotted in Fig. 14 B) and C) respectively. Similar to the flow velocity case, the values of invariants $\Psi_{A}$ and $\Psi_{B}$ range from 0 to 1. $\Psi_{A}$ corresponds to the concentration of reactant A and product C, exhibiting concentration in the upper half of the domain where reactant A is dominant, as per the problem setup. Similarly, $\Psi_{B}$ represents reactant B and product C, exhibiting concentration in the lower half, where reactant B is introduced in the problem setup.
@@
Finally, the concentration of reactants A and B along with product C are computed using the invariants $\Psi_{A}$ and $\Psi_{B}$ and presented in Fig. 15 . The mathematical calculation of the concentration is done using Eqs. (2.35)–(2.37). Reactant A exhibits its concentration at the top left boundary, gradually decreasing toward the right boundary. Similarly, reactant B is concentrated at the bottom left boundary and diminishes as it moves rightward. The formation of product C is observed at the interaction zone of reactants A and B, displaying a striking wavy concentration pattern. Product C begins forming at the left boundary and develops into a distinct plume extending toward the right boundary. Notably, the plume shifts slightly upward within the domain, which aligns with expectations, given that the coefficient $n_{B}$ is twice that of $n_{A}$, facilitating the upward movement of the plume.
@@ 6. CLOSURE
We have developed a physics-informed modeling framework for fast bimolecular reactions, incorporating tests for flow and reactive transport in porous media. In applications where fluid flow uncertainties and the risk of chemical contamination are concerns, PINNs—a novel machine learning framework—offers a powerful solution. Unlike traditional data-driven methods, PINNs is trained using the governing physics rather than relying on large datasets. For fast bimolecular reactions, traditional FEM and FVM solvers typically require fine spatial discretization, stabilization, or specialized treatments to resolve sharp mixing-limited fronts. In contrast, the proposed PINNs framework offers a mesh-free and flexible alternative that enforces governing equations and reaction constraints directly at collocation points. While not intended to outperform classical solvers in large-scale production settings, PINNs are particularly attractive for data-scarce problems, rapid prototyping, inverse modeling, and scenarios where repeated mesh generation and solver tuning become prohibitive.
This approach has broad applicability, ranging from critical minerals to subsurface energy systems such as geothermal energy and geological carbon storage, where fluid flow and chemical transport play critical roles. A key challenge in such subsurface applications is the scarcity of sufficient data. PINNs address this limitation by leveraging the system of PDEs and boundary conditions, enabling accurate modeling without the need for extensive experimental or field data. PINNs is trained by minimizing the total loss, which is a combination of individual losses calculated based on how well the solution at each collocation point satisfies the given PDE and boundary conditions. The system of equations varies depending on the nature of the problem—for instance, whether it involves flow or transport, as in this study. The overall computation time is problem-dependent, with the simplest configurations completing in approximately 2 minutes and the most refined cases in under an hour. During the development of the PINNs framework and solving the problems, several key observations were made, which are summarized below:
@@
Data supporting the findings of this study are available from the corresponding authors (KBN) on a reasonable request. The codes used in this paper will be made available and hosted at EMSL-Computing GitHub (https://github.com/EMSL-Computing/AI4Minerals) upon publication.